# Low-metallicity massive single stars with rotation

### v2

@@ 4 Available observations in the literature
Apart from the C iv 1550 Å line and the continuum photons, there are further observational features our results should be consistent with. To do such a comparison, we review the published spectral data of I Zw 18 focusing on individual emission lines/bumps. The figures below (Figs. 8–13) are reproduced from the literature for the sake of this discussion. Table 2 lists the papers, instruments and spectral ranges (optical or UV) mentioned here. As we will see, paying attention to the observing strategy of any given survey is key before drawing conclusions.
We remind that our best fit population predicts L 2.831037 erg s-1 (coming from twenty-something WO stars between 20-50 M⊙). This is actually close to what our newly computed PoWR model (with 131 M⊙ in the mid-pMS) predicts (2.131037 erg s-1). Therefore, to ease the following discussion, we will simply consider this mid-post main-sequence model (blue line in Fig. 4) representative for the combined spectrum of the population.
@@ 4.1 Optical emission lines/bumps
O VI 3818 Å. There are two observations available in the literature for this spectral region, Izotov et al. (1997) and Thuan & Izotov (2005). The first is reproduced here in Fig. 8 and the second, in Fig. 9. According to our model spectra in Fig. 4, this emission is predicted to be not a line but a huge double bump, ranging from 37803850 Å (see also Fig. 15 showing our line predictions). In the Izotov et al. (1997) spectra (Fig. 8), the flux calibration seems to be off at the position in question (it suddenly drops at the edge) making it hard to tell whether a broad component is there or not. The second paper, Thuan & Izotov (2005), did not report detection of the O VI 3818 Å line (see their Table 3), which is consistent with what one sees looking at Fig. 9. However, note that the Thuan-spectra also do not show the blue bump centered at 4650 which is so prominent in the Izotov-spectra. Thus, seeing other wide bumps of stellar origin should not be expected in the Thuan-data, either. We conclude that the currently available literature does not exclude our model predictions about this emission bump.
He II 4686 Å. This optical line is peculiar because it may have a stellar and a nebular component. The nebular one is narrow, the stellar (if there) broad, bump-like: our theory predicts a wide double bump between 4640-4700 Å (Fig. 4) which should be about as prominent as the C iv bump around 5808 Å. The Izotov et al. (1997) spectra therefore supports our theory: a bump is clearly there in Fig. 8 under the narrow nebular component. The same is true for the spectra presented in Kehrig et al. (2016), shown here in Fig. 10. As for Thuan & Izotov (2005), they measure an equivalent width of 2.5 Å (their Table 3) for the nebular line; whether there is a 60 Å-wide, almost completely cleared-out bump of stellar origin underneath it (Fig. 9), it is hard to say. We speculate that the lack of stellar emission bumps in the Thuan-data may be attributed to the extraction aperture (6”2”) not crossing over the star-clusters of interest, while the aperture of the Izotov-data (5”, centering on the North-Western region) includes these regions. Indeed, the position and size of the extraction apertures may play a key role in the detection of (or lack of) spectral features, as demonstrated by Kehrig et al. (2013, see their Fig. 11). We conclude that the existing literature does not exclude our model predictions about this line.
C IV 5808 Å. Only the Izotov et al. (1997) observation covers this region (Fig. 8). This measurement does show a bump at 5808, thus supporting our theory. Even the relative strength of it compared to the above discussed He II 4686 Å bump is tentatively consistent with our predictions in Fig. 4.
C IV 7724 Å. We did not found any spectra in the literature covering this wavelenght.
@@ 4.2 UV emission lines/bumps
He II 1640 Å. The appearance of this UV emission line (bump) is crucial, as it is typically used as a proxy for first stars (Pop-III). However, there are other ways to produce this line, as our models (massive Pop-II stars) attest. The bump is visible in the HST STIS spectra of Brown et al. (2002), as seen in Fig. 11, and its size is comparable to that of the C IV 1550 Å bump. Our theoretical populations nicely account for this. We predict line luminosities of lg L 37.58 for C iv (see Fig. 6) and lg L 37.35 for He ii (see Fig. 16) in the case of our mid-post-main-sequence WO model which dominates the population. As for the alternative population where a handful of more moderate-mass models dominate (cf. Sect. 3.1), the situation is similar: the values the WO star with Mini 59 M⊙ gives are lg L 36.82 for C iv and lg L 36.63 for He ii. Thus, this UV emission line supports our theory. We note however that when attempting to repeat the data reduction (the observations are publically available in the MAST archive555http://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html) we found it challenging to identify the prominent bumps reported by Brown et al. (2002). While reanalysing archival HST data is outside the scope of the current work, we encourage ongoing projects (e.g. S. Heap, private comm.) in this direction.
@@
O VI 1037 Å. This UV line is even stronger in our model spectra than C IV 1550 Å (see Figs. 4 and 16). Hence, we should look for it in the observations. Brown et al. (2002, Fig. 11) do not cover the position of O VI 1037 Å, unfortunately. Other works, such as Lecavelier des Etangs et al. (2004) do, although without covering C IV 1550 Å, as seen in Fig. 12. However, the observational strategy is again an important factor. Brown et al. (2002) took their UV spectrum as a slit crossing through the star-forming region in the North-Western cluster (see their Fig. 1 where they show their slits on the detailed map of I Zw 18, as well as our less detailed Fig. 11 where we show the same). Lecavelier des Etangs et al. (2004) applied another strategy: the FUSE aperture they used covers the full galaxy (30”x20”), as shown in Fig. 12. In this integrated spectrum, a stellar emission in O VI 1037 Å is not necessarily expected to be seen—the same way stellar emissions in the optical are not seen in the integrated spectrum presented by Kehrig et al. 2015b: the 4650 bump, otherwise identifiable in targeted observations (e.g. Izotov et al. 1997, see Fig 8) is completely cleared out when the data is summed over, as shown by Fig. 10 (reproduced from Fig. 3 of Kehrig et al. 2015b). Still, one might identify a P-Cygni profile around O VI1032 in the FUSE spectrum, just as we predict (cf. Fig. 16). We conclude that the available data in the literature does not exclude our model predictions about this line.
@@
C iv 1169 Å, C iii 1175 Å, O IV 1340 Å, O V 1371 Å. These lines/bumps are mentioned in Heap et al. (2015) who presents Hubble’s Cosmic Origins Spectrograph (COS) observations for them (their Fig. 5, reproduced here as Fig. 13). Since emission seems to be absent in their standard fit, they conclude that ,,the ultraviolet spectrum of [the North-Western region] shows no evidence of extremely hot, luminous stars”. The same conclusion was arrived at by Berg et al. (2022), after analysing the highest signal-to-noise-ratio UV spectrum of I Zw 18 available to date, taken by COS/HST (and incorporated in CLASSY). However, the observational strategy is again important here. As Fig. 13 shows, the COS field-of-view extends to a substantially large fraction of the North-Western region, as opposed to using slits crossing through the sources of interest (like the STIS data, see Fig. 14 for a direct comparison). As demonstrated by the integral-field-spectroscopy results of Kehrig et al. (2016, which we discussed in Fig. 10), emission bumps of stellar origin are only visible when the spectrum is extracted from the right spatial position when those massive stars (star-clusters) are physically located. When the entirety of the light is integrated over, the stellar emission clears out (Fig. 10, bottom). Therefore, we the fact that the COS data shows no stellar emission bumps does not necessarily exclude those stars being there. It may simply mean that the COS aperture is too large to detect them, and more targeted campaigns are needed. We conclude that these observations do not exclude our theory either.
O VI 2070 Å. We did not found any spectra in the literature covering this wavelenght.
@@ 4.3 On using MAST AstroView
While conducting this literature review, we have found a discrepancy in the AstroView tool of the MAST Archive. As shown by Figure 14, the two available backgrounds give strikingly different visuals at the position of I Zw 18 (RA 09:34:02, DEC +55:14:31). While the COS aperture’s position is the same, the DSS background (default option) is off by 4 arcsec compared to the Pan-STARRS background (alternative option in the AstroView settings). This can easily lead to confusion when interpreting the COS-data (or any other measurement).
According to the MAST Team (Y. Li, private comm.), the Pan-STARRS version is more accurate, although they have not confirmed this to be always true. They hope that these image products will be improved in the future, and they do not recommend using them for scientific analysis, only for preview. While for our present purposes, such previews are quite enough, this is a warning to the community that discrepancies of a few arcseconds can be expected when it comes to I Zw 18 and these images. For proper studies, coordinates from recent reference catalogs such as Gaia and Pan-STARRS1 should be used to confirm the accuracy of the galaxy’s—and especially its various regions’—celestial positions.
@@ 4.4 Summary of the literature review
As seen from this careful analysis of the literature—paying close attention to the observational strategies such as which apertures were used for which measurement—the currently available data are either supporting, and never excluding, our theoretical predictions. Even the Brown-data, which claims the presence of WC stars, do not exclude the possibility of these being, in fact, WO stars: the O iv 1340 Å and O V 1371 Å bumps are actually visible in their STIS spectra in Fig. 11. Our models predict these lines to be nowhere near as relevant as e.g. C IV 1550 Å or He II 4686 Å (Figs. 4 and 16), which is exactly what we see in the Brown-spectra. We conclude therefore that our suggestion about chemically-homogeneously evolved WO stars being responsible for the emission bumps so-far associated with WC stars is consistent with all available observations of I Zw 18.
@@ 5.1 Other possible explanations
The suggestion that nearly metal-free, hot massive stars (Pop III-like) are responsible for the He II emission in I Zw 18 (Kehrig et al. 2015b, a; Heap et al. 2015) is certainly appealing, but relies on the galaxy keeping (or obtaining) pockets of primordial gas that have just recently started to form massive stars for some unclear reason (cf. the review of Klessen & Glover 2023 claiming that Pop III star formation ended roughly at z 5). Our scenario follows more straightforwardly from the present day conditions in I Zw 18. Our stellar evolutionary models were computed with the measured composition of the very dwarf galaxy we study (following the gas composition reported in Lebouteiller et al. 2013), while our synthetic spectra were created based on these evolutionary models with the most state-of-the-art physics. The population synthesis is a rather general one, without extreme assumptions. So if metal-poor massive stars behave the way the models predict, then both Q and L are properly accounted for—without needing to postulate the presence of Pop III-like stars. Note that according to the seminal work of Yoon et al. (2012), Pop III stars also evolve chemically homogeneously when rotation is included, the same way as our metal-poor (i.e. massive Pop II) stars do.
@@
As for the scenario proposed by Péquignot (2008); Lebouteiller et al. (2017) and Schaerer et al. (2019) on X-ray binaries, we note that chemically-homogeneously evolving stars have been shown to lead to high-mass X-ray binaries in case they are born in a close binary system (Marchant et al. 2017). While Kehrig et al. (2021) refutes that high-mass X-rays binaries and/or diffuse soft X-ray photons would be the explanation for the high-ionizing features in I Zw 18, it may still be true in other galaxies. Hence, we suggest our scenario to be complementary to that of Schaerer et al. (2019), and welcome further research into the direction of combining these two possibilities (e.g. applying the methods developed in Sen et al. 2024 for predicting faint X-ray emissions).
@@
Indeed, while we only use single star models here, massive binaries are able to evolve chemically homogeneously too (e.g. de Mink et al. 2009; de Mink & Mandel 2016; Marchant et al. 2016). Alternatively, they may become analogous to our single, chemically homogeneous stars after losing their envelopes in a mass transfer phase (Götberg et al. 2017). Since double systems may be common in massive stellar populations (Sana et al. 2012; Hainich et al. 2018; Spencer et al. 2018), a more comprehensive picture of the contribution of chemically homogeneously evolving stars to the photoionization and carbon line emission in metal-poor galaxies will need to involve not only our fast rotating single stars (especially since there is ongoing debate on the effectiveness of rotationally induced mixing, e.g. Higgins & Vink 2019), but binaries as well.
@@
As for the model of Oskinova & Schaerer (2022) suggesting cluster-winds emitting X-rays, we note that this again can be nicely reconciled with our models. Indeed, our very evolutionary models have been applied in cluster-wind studies accounting for X-ray emission in Green Pea galaxies (Franeck et al. 2022). A similar study for I Zw 18 would be, therefore, quite possible, and we foresee a follow-up project combining the ionizing emission from massive stars and their shocked winds.
@@
Roy et al. (2025) recently suggested that classical WR stars with clumpy winds can produce hard photons in local and high-z galaxies. To test this scenario’s validity in I Zw 18, the C IV and other observed emission line luminosities that their models predict would need to be compared to observations. The same is true for the seminal work of Lecroq et al. (2024), in which they build a complex spectral synthesis tool including binary processes and investigate various emission lines (both nebular and stellar) in metal-poor stellar populations. We hope that by compiling existing UV and optical spectra from the literature on I Zw 18—one of the best studied dwarf galaxies out there—and by critically discussing this set of data in terms of its capacity to guide modelling efforts (Sect. 4), our work will assist future endeavors like these when testing their predictions directly against I Zw 18.

### v3

@@ 4 Available observations in the literature
Apart from the C iv 1550 Å line and the continuum photons, there are further observational features our results should be consistent with. To make such a comparison, we review here the published spectral data of I Zw 18 focusing on individual emission lines/bumps. The figures below (Figs. 8–13) are reproduced from the literature for the sake of this discussion. Table 2 lists the papers, instruments and spectral ranges (optical or UV) mentioned here. As we shall see, paying attention to the observing strategy of any given survey is key before drawing conclusions.
We remind the reader that our best fit population predicts L 2.831037 erg s-1 (coming from twenty-something WO stars between 20-50 M⊙). This is actually close to what our newly computed PoWR model (with 131 M⊙ in the mid-pMS) predicts (2.131037 erg s-1). Therefore, to ease the following discussion, we simply consider this mid-post-main-sequence model (blue line in Fig. 4) representative of the combined spectrum of the population.
@@ 4.1 Optical emission lines/bumps
O VI 3818 Å. There are two observations available in the literature for this spectral region, Izotov:1997 and Thuan:2005. The first is reproduced here in Fig. 8 and the second, in Fig. 9. According to our model spectra in Fig. 4, this emission is predicted to be not a line but a huge double bump, ranging from 37803850 Å (see also Fig. 15 showing our line predictions). In the Izotov:1997 spectra (Fig. 8), the flux calibration seems to be off at the position in question (it suddenly drops at the edge) making it hard to tell whether a broad component is there or not. The second paper, Thuan:2005, did not report detection of the O VI 3818 Å line (see their Table 3), which is consistent with what one sees looking at Fig. 9. However, the Thuan spectra also do not show the blue bump centred at 4650 that is so prominent in the Izotov spectra. Thus, seeing other wide bumps of stellar origin should not be expected in the Thuan data, either. We conclude that the currently available literature does not exclude our model predictions about this emission bump.
He II 4686 Å. This optical line is peculiar because it may have a stellar and a nebular component. The nebular one is narrow, the stellar (if there) broad, bump-like: our theory predicts a wide double bump between 4640-4700 Å (Fig. 4) that should be about as prominent as the C iv bump around 5808 Å. The Izotov:1997 spectra therefore supports our theory: a bump is clearly there in Fig. 8 under the narrow nebular component. The same is true for the spectra presented in Kehrig:2016, shown here in Fig. 10. As for Thuan:2005, they measure an equivalent width of 2.5 Å (their Table 3) for the nebular line; whether there is a 60 Å-wide, almost completely cleared-out bump of stellar origin underneath it (Fig. 9), it is hard to say. We speculate that the lack of stellar emission bumps in the Thuan data may be attributed to the extraction aperture (6”2”) not crossing over the star clusters of interest, while the aperture of the Izotov data (5”, centred on the north-western region) includes these regions. Indeed, the position and size of the extraction apertures may play a key role in the detection of (or lack of) spectral features, as demonstrated by Kehrig:2013. We conclude that the existing literature does not exclude our model predictions about this line.
C IV 5808 Å. Only the Izotov:1997 observation covers this region (Fig. 8). This measurement does show a bump at 5808, thus supporting our theory. Even the relative strength of it compared to the above discussed He II 4686 Å bump is tentatively consistent with our predictions in Fig. 4.
C IV 7724 Å. We did not found any spectra in the literature covering this wavelength.
@@ 4.2 UV emission lines/bumps
He II 1640 Å. The appearance of this UV emission line (bump) is crucial, as it is typically used as a proxy for first stars (Pop-III). However, there are other ways to produce this line, as our models (massive Pop-II stars) attest. The bump is visible in the HST STIS spectra of Brown:2002, as is seen in Fig. 11, and its size is comparable to that of the C IV 1550 Å bump. Our theoretical populations nicely account for this. We predict line luminosities of lg L 37.58 for C iv (see Fig. 6) and lg L 37.35 for He ii (see Fig. 16) in the case of our mid-post-main-sequence WO model that dominates the population. As for the alternative population where a select number of more moderate-mass models dominate (see Sect. 3.1), the situation is similar: the values the WO star with Mini 59 M⊙ gives are lg L 36.82 for C iv and lg L 36.63 for He ii. Thus, this UV emission line supports our theory. We note however that when attempting to repeat the data reduction (the observations are publicly available in the MAST archive666http://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html) we found it challenging to identify the prominent bumps reported by Brown:2002. While reanalysing archival HST data is outside the scope of the current work, we encourage ongoing projects (e.g. S. Heap, private comm.) in this direction.
@@
O VI 1037 Å. This UV line is even stronger in our model spectra than C IV 1550 Å (see Figs. 4 and 16). Hence, we should look for it in the observations. Brown:2002 do not cover the position of O VI 1037 Å, unfortunately. Other works, such as LecavelierdesEtangs:2004 do, although without covering C IV 1550 Å, as is seen in Fig. 12. However, the observational strategy is again an important factor. Brown:2002 took their UV spectrum as a slit crossing through the star-forming region in the north-western cluster (see their Fig. 1 where they show their slits on the detailed map of I Zw 18, as well as our less detailed Fig. 11 where we show the same). LecavelierdesEtangs:2004 applied another strategy: the FUSE aperture they used covers the full galaxy (30”x20”), as is shown in Fig. 12. In this integrated spectrum, a stellar emission in O VI 1037 Å is not necessarily expected to be seen – the same way stellar emissions in the optical are not seen in the integrated spectrum presented by Kehrig:2015: the 4650 bump, otherwise identifiable in targeted observations (e.g. Izotov:1997, see Fig 8) is completely cleared out when the data is summed over, as is shown by Fig. 10 (reproduced from Fig. 3 of Kehrig:2015). Still, one might identify a P-Cygni profile around O VI1032 in the FUSE spectrum, just as we predict (see Fig. 16). We conclude that the available data in the literature does not exclude our model predictions about this line.
@@
C iv 1169 Å, C iii 1175 Å, O IV 1340 Å, O V 1371 Å. These lines/bumps are mentioned in Heap:2015 who presents Hubble’s Cosmic Origins Spectrograph (COS) observations for them (their Fig. 5, reproduced here as Fig. 13). Since emission seems to be absent in their standard fit, they conclude that ,,the ultraviolet spectrum of [the north-western region] shows no evidence of extremely hot, luminous stars”. The same conclusion was arrived at by Berg:2022, after analysing the highest signal-to-noise-ratio UV spectrum of I Zw 18 available to date, taken by COS/HST (and incorporated in CLASSY). However, the observational strategy is again important here. As Fig. 13 shows, the COS field of view extends to a substantially large fraction of the north-western region, as opposed to using slits crossing through the sources of interest (similarly to the STIS data, see Fig. 14 for a direct comparison). As is demonstrated by the integral-field-spectroscopy results of Kehrig:2016, emission bumps of stellar origin are only visible when the spectrum is extracted from the right spatial position where those massive stars (star clusters) are physically located. When the entirety of the light is integrated over, the stellar emission clears out (Fig. 10, bottom). Therefore, we the fact that the COS data show no stellar emission bumps does not necessarily mean that those stars are not there. It may simply mean that the COS aperture is too large to detect them, and more targeted campaigns are needed. We conclude that these observations do not disprove our theory either.
O VI 2070 Å. We did not found any spectra in the literature covering this wavelength.
@@ 4.3 On using MAST AstroView
While conducting this literature review, we have found a discrepancy in the AstroView tool of the MAST Archive. As is shown by Figure 14, the two available backgrounds give strikingly different visuals at the position of I Zw 18 (RA 09:34:02, DEC +55:14:31). While the COS aperture’s position is the same, the DSS background (default option) is off by 4 arcsec compared to the Pan-STARRS background (alternative option in the AstroView settings). This can easily lead to confusion when interpreting the COS data (or any other measurement).
According to the MAST Team (Y. Li, private comm.), the Pan-STARRS version is more accurate, although they have not confirmed this to be always true. They hope that these image products will be improved in the future, and they do not recommend using them for scientific analysis, only for preview. While for our present purposes, such previews are quite enough, this is a warning to the community that discrepancies of a few arcseconds can be expected when it comes to I Zw 18 and these images. For proper studies, coordinates from recent reference catalogues such as Gaia and Pan-STARRS1 should be used to confirm the accuracy of the galaxy’s – and especially its various regions’ – celestial positions.
@@ 4.4 Summary of the literature review
As is seen from this careful analysis of the literature – paying close attention to the observational strategies such as which apertures were used for which measurement – the currently available data either support, or at least never exclude, our theoretical predictions. Even the Brown data, which claims the presence of WC stars, do not exclude the possibility of these being, in fact, WO stars: the O iv 1340 Å and O V 1371 Å bumps are actually visible in their STIS spectra in Fig. 11. Our models predict these lines to be nowhere near as relevant as for example C IV 1550 Å or He II 4686 Å (Figs. 4 and 16), which is exactly what we see in the Brown spectra. We conclude therefore that our suggestion about chemically homogeneously evolved WO stars being responsible for the emission bumps so far associated with WC stars is consistent with all available observations of I Zw 18.
@@ 5.1 Other possible explanations
The suggestion that nearly metal-free, hot massive stars (Pop III-like) are responsible for the He II emission in I Zw 18 (Kehrig:2015; Kehrig:2015b; Heap:2015) is certainly appealing, but relies on the galaxy keeping (or obtaining) pockets of primordial gas that have just recently started to form massive stars for some unclear reason (see the review of Klessen:2023 claiming that Pop III star formation ended roughly at z 5). Our scenario follows more straightforwardly from the present day conditions in I Zw 18. Our stellar evolutionary models were computed with the measured composition of the very dwarf galaxy we study (following the gas composition reported in Lebouteiller:2013), while our synthetic spectra were created based on these evolutionary models with the most state-of-the-art physics. The population synthesis is a rather general one, without extreme assumptions. So if metal-poor massive stars behave the way the models predict, then both Q and L are properly accounted for – without needing to postulate the presence of Pop III-like stars. Note that according to the seminal work of Yoon:2012, Pop III stars also evolve chemically homogeneously when rotation is included, the same way as our metal-poor (i.e. massive Pop II) stars do.
@@
As for the scenario proposed by Pequignot:2008; Lebouteiller:2017 and Schaerer:2019 on X-ray binaries, we note that chemically homogeneously evolving stars have been shown to lead to high-mass X-ray binaries in case they are born in a close binary system (Marchant:2017). While Kehrig:2021 refutes that high-mass X-rays binaries and/or diffuse soft X-ray photons would be the explanation for the high-ionizing features in I Zw 18, it may still be true in other galaxies. Hence, we suggest our scenario to be complementary to that of Schaerer:2019, and welcome further research into the direction of combining these two possibilities (e.g. applying the methods developed in Sen:2024 for predicting faint X-ray emissions).
@@
Indeed, while we only use single star models here, massive binaries are able to evolve chemically homogeneously too (e.g. deMink:2009b; deMink:2016; Marchant:2016). Alternatively, they may become analogous to our single, chemically homogeneous stars after losing their envelopes in a mass transfer phase (Goetberg:2017). Since double systems may be common in massive stellar populations (Sana:2012; Hainich:2018; Spencer:2018), a more comprehensive picture of the contribution of chemically homogeneously evolving stars to the photoionization and carbon line emission in metal-poor galaxies will need to involve not only our fast rotating single stars (especially since there is ongoing debate on the effectiveness of rotationally induced mixing, e.g. Higgins:2019), but binaries as well.
@@
As for the model of Oskinova:2022 suggesting cluster-winds emitting X-rays, we note that this again can be nicely reconciled with our models. Indeed, our very evolutionary models have been applied in cluster-wind studies accounting for X-ray emission in Green Pea galaxies (Franeck:2022). A similar study for I Zw 18 would be, therefore, quite possible, and we foresee a follow-up project combining the ionizing emission from massive stars and their shocked winds.
@@
Roy:2025 recently suggested that classical WR stars with clumpy winds can produce hard photons in local and high-z galaxies. To test this scenario’s validity in I Zw 18, the C IV and other observed emission line luminosities that their models predict would need to be compared to observations. The same is true for the seminal work of Lecroq:2024, in which they build a complex spectral synthesis tool including binary processes and investigate various emission lines (both nebular and stellar) in metal-poor stellar populations. We hope that by compiling existing UV and optical spectra from the literature on I Zw 18 – one of the best-studied dwarf galaxies out there – and by critically discussing this set of data in terms of its capacity to guide modelling efforts (Sect. 4), our work will assist future endeavours like these when testing their predictions directly against I Zw 18.